A nuclear reactor shuts down automatically if the temperature in the reactor differs from 1200 degrees Faherenhite by more than 30 degrees. For what temperatures will the reactor continue to operate?
Answer by stanbon(48568)

A nuclear reactor shuts down automatically if the temperature in the reactor differs from 1200 degrees Faherenhite by more than 30 degrees. For what temperatures will the reactor continue to operate?
-------------
Let "x" be a termperature at which the reactor will operate.
-------------
INEQUALITY:
|x-1200|<=30
-30<= x-1200 <=30
Add 1200 along the line to get:
1170 <= x <= 1230
==================
Cheers,
Stan H.